can you please explain to me how to do this and can you supply the answer so i can see how you got it thankx.
One number is two less than the second number. twice the first is four more than three times the second. find the positive difference between the product of the two numbers and their sum?
thankx for any help

Answer by ankor@dixie-net.com(12692)

Let x = "One number"
:
From the statement,"One number is two less than the second number." we can say:
Second number = (x+2)
:
Write an equation for statement (the word "is" usually means equal (=):
"twice the first is four more than three times the second."
2x = 4 + 3(x+2)
:
Multiply what's inside the bracket
2x = 4 + 3x + 6
2x - 3x = 6 + 4; subtract 3x from both sides
-x = 10
x = -10; multiplied the equation by -1
:
The second number = x + 2; therefore: -10 + 2 = -8
:
find the positive difference between the product of the two numbers and their sum?
Find the product: -10 * -8 = +80
Find the sum: -10 + (-8) = -18
----------------------------------use the rules of subtraction to find the difference
The positive difference = +98
:
:
Could you see how we did this?